### v1

# It’s Morphing Time: Unleashing the Potential of Multiple LLMs via Multi-objective Optimization

Abstract
In this paper, we introduce a novel approach for large language model merging via black-box multi-objective optimization algorithms. The goal of model merging is to combine multiple models, each excelling in different tasks, into a single model that outperforms any of the individual source models. However, model merging faces two significant challenges: First, existing methods rely heavily on human intuition and customized strategies. Second, parameter conflicts often arise during merging, and while methods like DARE [1] can alleviate this issue, they tend to stochastically drop parameters, risking the loss of important delta parameters. To address these challenges, we propose the MM-MO method, which automates the search for optimal merging configurations using multi-objective optimization algorithms, eliminating the need for human intuition. During the configuration searching process, we use estimated performance across multiple diverse tasks as optimization objectives in order to alleviate the parameter conflicting between different source models without losing crucial delta parameters. We conducted comparative experiments with other mainstream model merging methods, demonstrating that our method consistently outperforms them. Moreover, our experiments reveal that even task types not explicitly targeted as optimization objectives show performance improvements, indicating that our method enhances the overall potential of the model rather than merely overfitting to specific task types. This approach provides a significant advancement in model merging techniques, offering a robust and plug-and-play solution for integrating diverse models into a unified, high-performing model.
1 Introduction
Large language models (LLMs) have shown great performance on tasks in various domains such as natural language processing [2], computer vision [3] etc. With the great effort of world-wide contributors in community, a large number of general-purpose pre-trained and task-specific fine-tuned language models have been proposed and made publicly available. However, LLM pre-training or fine-tuning is non-trivial and requires a lot of effort and financial budget. Recently, model merging (MM) has attracted many researchers’ attention. By combining multiple LLMs into a single model with better performance and adaptability on more tasks, MM offers a novel cost-efficient way of obtaining new powerful language models without performing additional training [4, 5], just like the Power Rangers merge their Zords together to form a mighty Megazord [6]. Ideally, MM is supposed to inherit and amplify the strengths from its source models while ignoring their weaknesses. Therefore, the obtained model will be able to tackle the union set of all the tasks where the source models are pre-trained/fine-tuned with better performance. Yet this is achieved without training, which saves a large amount of calculation/financial budget. With the help of open-source toolkits such as mergekit [7, 8], MM has become popular for LLM developing and shown great potential on the Open LLM Leaderboard [9]
However, model merging requires the model maker to have profound knowledge or intuition. automatically discover more capable MM recipes is still in its infancy. To the best of our knowledge, The only related work is [4], where Akiba et al. proposed to use evolutionary algorithms (EAs) to generate powerful model merging recipes which operates in both parameter space and data flow space. However, the approach of using diverse task scores and optimizing the parameter space to explore the comprehensive potential of the final model has been largely overlooked.
MM is similar to ensemble learning in the sense that they both try to produce one model from multiple source models. In ensemble learning, It is generally believed that diversity in an ensemble could be beneficial for improving its performance [10, 11, 12]. One could infer that maintaining diversity during MM may also result in powerful LLMs. Having this in mind, we design a Model Merging approach via multi-objective optimization which take into consideration multiple task-specific performance metrics simultaneously.
In this work, we proposed to leveraging multi-objective optimization for automatic model merging. Specifically, we reformulate the problem of merging multiple models as a multi-objective optimization problem. This automated model merging approach offers several advantages over retraining a new model from scratch: Every model has its unique strengths and weaknesses based on its training dataset and architecture. Merging complementary models into one hyper-efficient system results in a performance powerhouse that leverages the best qualities of each model. The contributions of this work can be summarized as follows:
(1) Automated Model Merging: We introduce a novel approach to model merging that automates the search for optimal merging configurations using multi-objective black-box optimization algorithms, eliminating the reliance on human intuition and customized strategies.
(2) Enhanced Model Potential: Our experiments demonstrate that the proposed method not only preserves individual capabilities while mitigating parameter interference but also improves performance on problem types not explicitly targeted as optimization objectives. This indicates that our method enhances the overall potential of the model, providing a robust and versatile solution for integrating diverse task-specific models into a high-performing, cohesive model.
(3) Comprehensive Evaluation: We validate the effectiveness of our MM-MO method through extensive evaluations on three different types of natural language tasks. The results show that the models merged using our proposed approach achieve superior performance compared to existing methods, highlighting the advantages of our method.
2 Background
2.1 Model Merging
Generally speaking, there are four kinds of model merging (MM) methods.
One common approach is simply averaging the weights of different models, which has shown promising performance. Model soups [13] achieved notable performance improvements on computer vision tasks. Task arithmetic [14] merges large language models (LLMs) by performing arithmetic operations on task vectors (also known as delta parameters), which represent the differences between the weights of fine-tuned LLMs and the original pre-trained model. Fisher-Weighted Averaging [15] is based on the Laplace approximation, where each model’s posterior is approximated as a Gaussian distribution, with its precision matrix corresponding to its Fisher information.
Spherical linear interpolation (SLERP) [16] can achieve smooth interpolation between two vectors and maintain directional consistency during model merging by calculating the angle between two vectors and determining the interpolation vector based on the interpolation factor and angle.
RegMean [17] merges different models by minimizing the prediction differences between the merged model and the source models.
To address the interference between parameters of different models caused by redundant parameter values and disagreements on the sign of parameters across models, TIES-Merging [18] introduces three novel steps when merging models: (1) resetting parameters with small delta parameters, (2) resolving sign conflicts, and (3) merging only the consistent parameters.
DARE [19] performs model merging by sparsifying delta parameters of source supervised fine-tuned (SFT) models with Drop and Re-Scale operations before actually merging them.
Recently, Daheim et al. [5] connected the inaccuracy of weighted-averaging to mismatches in the gradients and proposed a new uncertainty-based scheme to improve performance by reducing this mismatch.
Our approach builds upon the combination of TIES-Merging and DARE, leveraging their strengths to create a robust model merging strategy. By integrating the conflict resolution techniques of TIES-Merging with the parameter sparsification methods of DARE, we aim to automate the search for optimal model merging configurations, ensuring that the final merged model maintains high performance without the degradation typically associated with parameter conflicts.
2.2 Relationship between MM and Multi-Task Learning and Ensemble Learning
Multitask learning (MTL) is a learning mechanism that improves learning efficiency and performance by learning multiple tasks simultaneously. In deep learning, MTL refers to training a neural network to perform multiple tasks by sharing some of the network’s layers and parameters across tasks. MM is similar to MTL since they both try to obtain a machine learning model that can handle multiple tasks when one aim to produce a LLM for different tasks with MM.
Ensemble Learning (EL) Ensemble learning is a machine learning technique that aggregates two or more learners (e.g. regression models, neural networks) in order to produce better predictions. MM is similar to EL since they both try to obtain a machine learning model with multiple source models.
Roughly speaking, MM resembles MTL in final goal (ability on multiple tasks) and EL in input with multiple source models. Other than the above-mentioned points, MM differs from both MTL and EL in that it involves no further training or fine-tuning.
2.3 Multi-objective Optimization
Without loss of generality, a multi-objective optimization problem can be stated as follows:
$&\min\textbf{ }\boldsymbol{f}(\boldsymbol{x})=(f_{1}(\boldsymbol{x}),f_{2}(\boldsymbol{x}),\ldots,f_{M}(\boldsymbol{x}))^{T}\\ &\textrm{ }s.t.\textrm{ }\boldsymbol{x}\in\mathcal{X}\end{aligned},$ (1)
where $\boldsymbol{x}$ $=$ $(x_{1},x_{2},\ldots,x_{d})$ is the decision vector, $\boldsymbol{f}(\cdot)$: $\mathcal{X}\rightarrow\mathcal{T}$ is $m$ objective functions, $\mathcal{X}$ and $\mathcal{T}$ denote the (nonempty) decision space and the objective space, respectively.
In order to compare the quality of solutions of an MOP, the concept of Pareto dominance is introduced:
Definition 2.1 (Pareto dominance [20])
Given two solutions $\boldsymbol{a}$, $\boldsymbol{b}\in$ $\mathcal{X}$, $\boldsymbol{a}$ is said to dominate $\boldsymbol{b}$ (denoted as $\boldsymbol{a}\prec\boldsymbol{b}$) if and only if $\forall i\in\{1,2,...,m\},f_{i}(\boldsymbol{a})\leq f_{i}(\boldsymbol{b})$ and $\exists j\in\{1,2,...,m\},f_{j}(\boldsymbol{a})<f_{j}(\boldsymbol{b})$. A solution $\boldsymbol{a}^{\ast}\in\mathcal{X}$ is Pareto optimal if no other solution $\boldsymbol{a}\in\mathcal{X}$ can dominate it. The solution set consisting of all the Pareto optimal solutions is called the Pareto set (PS): $PS$$=$$\{\boldsymbol{a}\in\mathcal{X}|\forall\boldsymbol{b}\in\mathcal{X},\boldsymbol{b}\not\prec\boldsymbol{a}\}$ and the corresponding objective vector set of the PS is the Pareto front (PF).
Multi-objective Optimization (MO) focuses on approximating the PS, targeting at a solution set with good convergence and diversity in the objective space. In this work, we utilize multi-objective optimization to address the complex challenge of model merging. By considering multiple objectives simultaneously, MO enables us to balance various performance metrics, such as different task-specific accuracies. This approach allows us to optimize the trade-offs between different performance criteria, leading to a more robust and effective merged model. Through this methodology, we aim to achieve a unified model that not only excels in specific tasks but also retains a broad and diverse set of capabilities from the individual models. This is particularly important in the context of large language models, where different tasks may require different strengths and features from the base models. By leveraging MO, our method systematically explores the parameter space to identify optimal configurations, ensuring that the final model exhibits both high performance and comprehensive potential across various tasks.
3 Methodology
3.1 Overview
In this section, we provide an overview of our proposed automated model merging approach, MM-MO. Our approach is designed to merge multiple pre-trained models into a single, unified model that retains the strengths of each individual model and surpasses the performance of any single model. Our automated approach aims at searching for the optimal merging configuration is discovered without relying on human intuition.
Figure 1: An illustration of automated model merging with multi-objective optimization (MM-MO).
In Figure 1, The dashed lines indicate that the model exhibits relatively superior performance on the corresponding tasks. The final output is the merged model, which integrates the strengths of the individual models and demonstrates superior performance on a wide range of tasks. Our method starts with a collection of source models, each fine-tuned on different tasks. The objective is to merge these diverse models into a single model that not only performs well on the individual tasks but also demonstrates enhanced overall capability across a range of tasks.
The merging process leverages a collection of tasks to guide the optimization. Each source model is evaluated across this task collection to determine its strengths and weaknesses. The task collection includes a variety of tasks, ensuring a comprehensive evaluation of model performance. To improve the ability of the final merged model, we utilize multi-objective black-box optimization. This process involves balancing different performance metrics, such as task-specific accuracy values, while preserving crucial delta parameters. The goal is to achieve a merged model that exhibits both high performance and versatility.
3.2 Model Merging via Multi-objective Optimization
In this subsection, we detail the multi-objective optimization process employed for model merging. Our method leverages parallel multi-objective bayesian optimization (qEHVI) [21]. This technique is an effective black-box optimization method designed to handle multiple objectives simultaneously, making it ideal for our model merging needs.
Our approach builds upon TIES-Merging with DARE [18, 1], utilizing qEHVI to search for the optimal model merging configuration. The integration of qEHVI allows for an automated and systematic exploration of the parameter space, identifying the best configurations that balance various performance metrics. Here are the detailed steps of our methodology.
  • 1)
    Initialization: We start with a collection of pre-trained models, each fine-tuned on different tasks. These models and their configurations serve as the initial candidates for merging.
  • 2)
    Objective Definition: We define multiple objectives that reflect the performance metrics we aim to optimize. In our case, these objectives include task-specific accuracies such as C-EVAL [22] and GSM8K [23] accuracy. These metrics are crucial as they ensure the merged model performs well across different types of tasks.
  • 3)
    Surrogate Model Training: qEHVI leverages surrogate models to approximate the objective functions. We train these surrogate models on the initial set of model configurations and their corresponding performance metrics. This step involves creating a probabilistic model that predicts the performance of unseen configurations.
  • 4)
    Acquisition Function Optimization: The key component of qEHVI is the Expected Hypervolume Improvement (EHVI) acquisition function. This function guides the search by quantifying the expected improvement in the objective space. We optimize this acquisition function to select the most promising model configurations to evaluate next. Specifically, the hypervolume (HV) is a measure of the space dominated by a set of solutions in the objective space, bounded from below by a reference point. The HV indicator is defined as follows:
    $\textbf{HV}(S)=\Lambda(\{q\in\mathbb{R}^{d}|\exists p\in S:p\leq q\;\textbf{and}\;q\leq r\}),$ (2)
    where $S$ signifies a solution set, $r$ is identified as a reference vector set, and $\Lambda(\cdot)$ denotes the Lebesgue measure. Hypervolume improvement (HVI) measures the additional hypervolume that a new point contributes beyond the existing Pareto frontier.
    In black-box optimization, exact function values at unobserved points are unknown, making direct HVI computation infeasible. Instead, the EHVI acquisition function uses the surrogate model’s posterior distribution over the function values to compute the expected improvement. This is typically done using Monte Carlo (MC) integration to estimate EHVI [24], particularly when dealing with multiple candidates in parallel. The EHVI guides the search by identifying new points that are expected to provide the most significant improvement in the objective space. The more general parallel variant utilizing MC integration is given as follows:
    $\alpha_{q\text{EHVI}}(\mathcal{X}_{\text{cand}}|\mathcal{P})\approx\hat{\alpha}_{q\text{EHVI}}(\mathcal{X}_{\text{cand}}|\mathcal{P})=\frac{1}{N}\sum_{t=1}^{N}\text{HVI}(\tilde{\boldsymbol{f}}_{t}(\mathcal{X}_{\text{cand}})|\mathcal{P}),$ (3)
    where $\tilde{f}_{t}\sim p(\boldsymbol{f}|\mathcal{D})$ for $t=1,...,N$ and $\mathcal{X}_{\mathrm{cand}}=\{x_{i}\}_{i=1}^{q}$ [21].
  • 5)
    Parallel Evaluation: Unlike traditional methods, qEHVI supports parallel evaluations, which significantly speeds up the optimization process. In each iteration, we select a batch of model configurations and evaluate their performance on the defined objectives (e.g., C-EVAL and GSM8K accuracy).
  • 6)
    Updating the Surrogate Model: After evaluating the selected configurations, we update the surrogate models with the new data. This iterative process continues, refining the surrogate models and improving the acquisition function’s accuracy in identifying high-potential configurations.
  • 7)
    Convergence and Final Selection: The optimization process continues until convergence criteria are met, such as a maximum number of iterations or negligible improvement in the objective space. The final model merging configuration is selected based on the best trade-offs between the multiple performance metrics.
By using multi-objective Bayesian optimization method, our method systematically navigates the trade-offs between different task-specific accuracy, ensuring that the final merged model maintains high performance across various tasks. The iterative nature of this process allows for continuous improvement and fine-tuning, leading to a robust and high-performing unified model. This integration of qEHVI with TIES-Merging and DARE provides a powerful and automated solution for model merging, optimizing the balance between multiple performance criteria without relying on human intuition. This automated search for the optimal merging configuration ensures that the final model configuration surpasses the performance of any individual source model.
4 Experimental Study
4.1 Setup
Source Models
To develop a model with strong comprehensive abilities across various disciplines, we applied our proposed model merging method to a set of source models, including Qwen1.5-7B-Chat [25], Liberated-Qwen1.5-7B (https://huggingface.co/abacusai/Liberated-Qwen1.5-7B), and firefly-qwen1.5-en-7B (https://huggingface.co/YeungNLP/firefly-qwen1.5-en-7b). All these models were fine-tuned from Qwen1.5-7B [25].
Datasets
We chose C-EVAL [22], a Multi-Level Multi-Discipline Chinese Evaluation Suite, to test the comprehensive abilities of our models across different disciplines. Additionally, we employed a variety of datasets to assess the models’ performance across diverse task scenarios. Specifically, we used GSM8K [23] to evaluate the models’ mathematical reasoning abilities and HumanEval [26] to test their code generation capabilities. Moreover, we used a different dataset for BO search to avoid overfitting the test set. Specifically, we use the valid splits of C-EVAL that were not included in the C-EVAL test set, and we use 1% of the GSM8K train set sampled randomly to ensure diversity and to prevent overfitting.
Evaluation Metrics
C-EVAL: This dataset is used for tasks that involve classification, such as true/false questions and multiple-choice questions. Since these types of tasks require selecting the correct option from a set of possible answers, we use the accuracy metric to evaluate performance. Specifically, we calculate the 5-shot accuracy, which measures the model’s accuracy when given five examples for context before making predictions. This approach helps in assessing the model’s ability to learn from a few examples and generalize to similar tasks. GSM8K: This dataset is designed for mathematical problem-solving tasks, which often require generating accurate and precise solutions from scratch. Since these tasks do not have multiple-choice answers but rather require an exact answer, we evaluate performance using zero-shot accuracy. This means we assess the model’s ability to correctly solve problems without any prior examples or fine-tuning, reflecting its inherent capability to handle complex, unseen mathematical tasks. HumanEval: This dataset is used for evaluating code generation tasks. The nature of these tasks requires generating functional and correct code based on a given problem description. We use the pass@1 metric to evaluate performance, which measures the proportion of problems for which the first generated solution passes all test cases. This metric is crucial as it directly assesses the model’s ability to produce correct and executable code on the first attempt, which is a stringent but realistic measure of practical utility in coding applications.
4.2 Results and Analysis
4.2.1 General Performance
Table 1: Performance Comparison of Different Merging Methods and Single Models
Merging Methods Source Models C-EVAL GSM8K Human Eval
Single Model 1 Qwen/Qwen1.5-7B-Chat 67.4 60.27 46.95
Single Model 2 abacusai/Liberated-Qwen1.5-7B 69.9 32.60 28.04
Single Model 3 YeungNLP/firefly-qwen1.5-en-7b 70.1 48.36 20.12
Task Arithmetic Single Model 1 + 2 + 3 70.2 58.83 50.60
TIES Single Model 1 + 2 + 3 69.9 54.20 42.07
DARE + TIES Single Model 1 + 2 + 3 70.0 51.32 50.60
MM-MO (Ours) Single Model 1 + 2 + 3 71.4 66.56 56.09
Table 1 summarizes the performance of different merging methods and single models on the C-EVAL, GSM8K, and HumanEval tasks. Single Model 1 (Qwen/Qwen1.5-7B-Chat) achieved scores of 67.4, 60.27, and 46.95 on C-EVAL, GSM8K, and HumanEval, respectively, demonstrating relatively balanced performance. Single Model 2 (abacusai/Liberated-Qwen1.5-7B) scored higher on C-EVAL with 69.9 but showed weaker performance on GSM8K and HumanEval, with scores of 32.60 and 28.04, respectively. Single Model 3 (YeungNLP/firefly-qwen1.5-en-7b) performed reasonably well on C-EVAL and GSM8K, scoring 70.1 and 48.36, respectively, but had the lowest score on HumanEval, at 20.12.
Different merging methods significantly impact model performance. The Task Arithmetic method achieved scores of 70.2, 58.83, and 50.60 on C-EVAL, GSM8K, and HumanEval, respectively, indicating consistent improvements across all tasks. The TIES method showed a slight decrease in performance on GSM8K with a score of 54.20, but maintained good performance on C-EVAL and HumanEval, with scores of 69.9 and 42.07, respectively. The DARE + TIES method had consistent scores with the Task Arithmetic method on HumanEval, both achieving 50.60, but showed slightly lower performance on C-EVAL and GSM8K, with scores of 70.0 and 51.32, respectively. Our method (MM-MO) is based on DARE + TIES and incorporates a multi-objective optimization algorithm designed based on the performance of merged models on the training sets of the C-EVAL and GSM8K datasets to search for the optimal merging configuration. Consequently, it achieved scores of 71.4, 66.56, and 56.09 on C-EVAL, GSM8K, and HumanEval, respectively, demonstrating superior performance across all tasks. Notably, our method exhibited significant improvements on GSM8K and HumanEval compared to single models and other merging methods.
In summary, although the individual models show varying performance across different tasks, effective model merging strategies can substantially enhance overall performance. Our MM-MO method consistently achieved the best results in all evaluated tasks, validating the significant potential of merged models in tackling complex tasks.
4.2.2 Case Study
Table 2: Case Study of MM-MO. This case study shows the responses of different merged models to a problem that demands mathematical reasoning.
Problem
Instruction Mark took a test yesterday that consisted of 75 questions. He completed the test at a rate of 5 questions per hour. Today, he took another test of 100 questions at the same rate. If Mark had 8 hours to complete the first test and 6 hours to complete the second one, how many questions did he leave incomplete?
Ground Truth Answer 105
Note Questions Mark completed in yesterday’s test: 5*8=40. Questions Mark completed in today’s test: 5*6=30. Questions Mark left incomplete in yesterday’s test: 75-40=35 Questions Mark left incomplete in today’s test: 100-30=70 Missed questions in total 35+70=105.
Answers
MM-MO (Ours) To find out how many questions Mark left incomplete, we need to calculate the number of questions he completed in each test and then subtract that from the total number of questions in each test. For the first test: - Mark completed 5 questions per hour for 8 hours, so he completed 5 * 8 = 40 questions. - The total number of questions in the first test is 75, so he left 75 - 40 = 35 questions incomplete. For the second test: - Mark completed 5 questions per hour for 6 hours, so he completed 5 * 6 = 30 questions. - The total number of questions in the second test is 100, so he left 100 - 30 = 70 questions incomplete. Now, to find the total number of questions left incomplete, we add the number of questions left incomplete in each test: 35 (first test) + 70 (second test) = 105 questions The answer is: 105 questions.
TIES To find out how many questions Mark left incomplete, we need to calculate the total number of questions he completed in each test and then subtract that from the total number of questions in each test. For the first test: - Number of questions completed: 75 questions / 5 questions per hour = 15 hours - Number of questions left incomplete: 8 hours - 15 hours = -7 hours (This is not possible, so there are no questions left incomplete for the first test) For the second test: - Number of questions completed: 100 questions / 5 questions per hour = 20 hours - Number of questions left incomplete: 6 hours - 20 hours = -14 hours (This is not possible, so there are no questions left incomplete for the second test)
DARE-TIES You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day!
Table 2 provides a comparative analysis of responses to a mathematical reasoning problem by three models, illustrating the remarkable improvement in performance through the introduction of a multi-objective optimization algorithm in our MM-MO method. By employing multi-objective optimization methods to search for the optimal configuration for model merging, we not only enhanced our model’s performance in the multi-disciplinary evaluation test (C-EVAL, Table 1) and code generation (Human Eval, Table 1) but also further uncovered its mathematical reasoning capabilities.
Specifically, in this problem, only our merged model gives the correct answer. Our model (merged by MM-MO method) understood the question well and followed the step by step instructions to give the final correct answer. However, the model merged by TIES method, while also following the step by step instructions, does not properly understand the problem and give a mathematical reasoning process that meets our requirements. Finally, the model merged by DARE-TIES method fails entirely, providing irrelevant and repetitive responses, despite DARE-TIES utilizing a dropout-like method to avoid parameter conflicts during model merging, which generally results in better performance compared to TIES, it can suffer from significant performance drops on specific problems if the model merging configuration is not correctly set. This issue is evident in the provided case study, where DARE-TIES fails to deliver a coherent response. This highlights the importance of correctly configuring model merges to retain crucial parameters, ensuring consistent performance across diverse problem sets. The MM-MO model’s success demonstrates the effectiveness of our model merging strategy in overcoming these challenges and providing accurate solutions.
4.2.3 Ablation Study
Table 3: Ablation study: (1) DARE-TIES without MM-MO (2) MM-MO, consider only the accuracy of C-EVAL (3) MM-MO, consider both the accuracies of C-EVAL and GSM8K (Default setting).
Merging Methods Source Models C-EVAL GSM8K Human Eval
DARE + TIES Single Model 1 + 2 + 3 70.0 51.32 50.60
MM-MO
(1 obj: C-EVAL) Single Model 1 + 2 + 3 71.5 66.18 51.82
MM-MO
(2 objs: C-EVAL & GSM8K) Single Model 1 + 2 + 3 71.4 66.56 56.09
In this ablation study, we compared the impact of different model merging methods on model performance, specifically comparing DARE + TIES and two types of MM-MO (single-objective and multi-objective) methods. The table presents the performance of these three merging methods across three evaluation tasks (C-EVAL, GSM8K, and Human Eval).
First, using the default configuration (average weights), the DARE + TIES method serves as the baseline, achieving scores of 70.0, 51.32, and 50.60 on C-EVAL, GSM8K, and Human Eval, respectively. Next, we introduced the proposed MM-MO method. Initially, we considered the MM-MO method with a single objective focused solely on C-EVAL. This method increased the score on C-EVAL to 71.5, indicating that our black-box optimization strategy in model merging can enhance performance on specific tasks. Moreover, since C-EVAL is a comprehensive academic assessment test set, improving the accuracy of the merged model on C-EVAL also brought enhancements in mathematical reasoning and code generation capabilities. Specifically, the model obtained through the single-objective MM-MO merging strategy achieved scores of 66.18 and 51.82 on GSM8K and Human Eval, respectively, outperforming our baseline.
Finally, we employed the MM-MO method considering both C-EVAL and GSM8K as objectives. This approach yielded scores of 71.4, 66.56, and 56.09 on C-EVAL, GSM8K, and Human Eval, respectively. Compared to the single-objective MM-MO method focusing only on C-EVAL, this approach achieved comparable results on C-EVAL and a further notable improvement in the accuracy on GSM8K. This infers that the mathematical reasoning (GSM8K) capability of LLM may contribute to code generation (Human Eval) ability. Additionally, although Human Eval was not the best, the increased number of optimization objectives enhanced the model’s overall capabilities, significantly improving its code generation ability, with the score increased to 56.09. This demonstrates that considering multiple objectives can achieve better trade-off and overall performance across different tasks.
In conclusion, the proposed MM-MO method, by incorporating multi-objective black-box optimization, effectively enhances model performance across various tasks, underscoring the importance and advantages of multi-objective optimization in model merging configurations.
5 Limitation
Our experiments were constrained by the available computational resources, limiting the application of our method to the merging of models with up to 7 billion parameters. While our approach demonstrated promising results within this parameter range, the effectiveness and scalability of our method for models with significantly larger parameter counts remain untested. Future research is necessary to evaluate the performance and feasibility of our multi-objective optimization method for merging models with higher parameter volumes. This exploration will be crucial to understand the full potential and limitations of our approach in handling more extensive and complex models.
Additionally, our model merging method currently requires that the source models be homologous, meaning they must be fine-tuned versions of the same pre-trained model. This homogeneity constraint ensures compatibility during the merging process. However, it also limits the applicability of our method to scenarios where source models are derived from different pre-trained models. Future work should explore the feasibility of extending our approach to handle heterogeneous source models, broadening the scope and utility of our model merging technique.
6 Conclusion
This paper proposes a novel approach for large language model merging using multi-objective black-box optimization algorithms, addressing two major challenges: the reliance on human intuition for merging configurations and the potential loss of crucial parameters with DARE strategies. By employing qEHVI, our method automates the search for optimal merging configurations, ensuring that the merged model maintains high performance without the typical parameter conflict issues.
Model merging offers significant advantages over training models from scratch, including enhanced knowledge integration, performance optimization, and cost efficiency. Our approach effectively combines specialized knowledge from multiple pre-trained models, creating a unified model that leverages the best qualities of each, resulting in superior overall performance. Additionally, merging models is more resource-efficient, requiring less energy and computational power.
Our contributions include automating the model merging process, enhancing model potential by preserving individual capabilities and improving performance on diverse tasks, and validating our method through extensive evaluations on various natural language tasks. The results demonstrate that our approach achieves superior performance compared to existing methods, offering a robust, high-performing solution for integrating diverse task-specific models.
In summary, our multi-objective optimization method provides an automated, efficient, and powerful solution for model merging, addressing key challenges and resulting in robust, high-performing models that leverage the strengths of multiple pre-trained models.