### v2

# It’s Morphing Time: Unleashing the Potential of Multiple LLMs via Multi-objective Optimization

Abstract
In this paper, we introduce a novel approach for large language model merging via black-box multi-objective optimization algorithms. The goal of model merging is to combine multiple models, each excelling in different tasks, into a single model that outperforms any of the individual source models. However, model merging faces two significant challenges: First, existing methods rely heavily on human intuition and customized strategies to tackle multiple tasks. Second, it’s difficult to search for the great model merging configuration in limited evaluations. To address these challenges, we propose a multi-objective optimization based model merging method named MM-MO. The proposed method can automatically search merging configurations for multiple tasks with multi-objective optimization algorithms. Moreover, to obtain high-quality model merging configurations within a limited number of evaluation iterations, we have made several improvements to multi-objective Bayesian optimization specifically for model merging scenarios. First, we introduced a weak-to-strong method to improve the acquisition strategy. Second, we employed Fisher information to select configurations, further increasing the chances of discovering superior model merging configurations. Third, we designed a sparsity metric as an additional optimization objective to enhance the model’s generalization performance across different tasks. We conducted comprehensive experiments with other mainstream model merging methods, demonstrating that our method consistently outperforms them. Moreover, performance improvements are observed even on the tasks not explicitly targeted as optimization objectives, indicating that our method enhances the overall potential of the model. The proposed method provides a significant advancement in model merging techniques, offering a robust and efficient solution for integrating diverse models into a unified, high-performing model without additional training data.
1 Introduction
Large language models (LLMs) have shown great performance on tasks in various domains such as natural language processing [1], computer vision [2] etc. With the great effort of world-wide contributors in community, a large number of general-purpose pre-trained and task-specific fine-tuned language models have been proposed and made publicly available. However, LLM pre-training or fine-tuning is non-trivial and requires a lot of effort and financial budget. Recently, model merging (MM) has attracted many researchers’ attention. By combining multiple LLMs into a single model with better performance and adaptability on more tasks, MM offers a novel cost-efficient way of obtaining new powerful language models without performing additional training [3, 4], just like the Power Rangers merge their Zords together to form a mighty Megazord [5]. Ideally, MM is supposed to inherit and amplify the strengths from its source models while ignoring their weaknesses. Therefore, the obtained model will be able to tackle the union set of all the tasks where the source models are pre-trained/fine-tuned with better performance. Yet this is achieved without training, which saves a large amount of calculation/financial budget. With the help of open-source toolkits such as mergekit [6, 7], MM has become popular for LLM developing and shown great potential on the Open LLM Leaderboard [8].
However, model merging requires the model maker to have profound knowledge or intuition. Automatically discovering more capable MM recipes is still in its infancy. To the best of our knowledge, The most related work is [3], where Akiba et al. proposed to use evolutionary algorithms (EAs) to generate powerful model merging recipes which operates in both parameter space and data flow space. However, the approach of using diverse task scores and optimizing the parameter space to explore the comprehensive potential of the final model has been largely overlooked.
MM is similar to ensemble learning in the sense that they both try to produce one model from multiple source models. In ensemble learning, It is generally believed that diversity in an ensemble could be beneficial for improving its performance [9, 10, 11]. One could infer that maintaining diversity during MM may also result in powerful LLMs. Having this in mind, we design a Model Merging approach via multi-objective optimization which take into consideration multiple task-specific performance metrics simultaneously.
In this work, we proposed to leverage multi-objective optimization for automatic model merging. Specifically, we formulate the problem of merging multiple models as a multi-objective optimization problem. This automated model merging approach is able to obtain a performance powerhouse that leverages the best qualities of each model, inheriting their respective strengths of all source models. Furthermore, we validate the effectiveness of our MM-MO method through extensive evaluations across a diverse range of LLM evaluation tasks. The contributions of this work can be summarized as follows:
(1) Automated Model Merging via multi-objective optimization: We introduce a novel approach to model merging that automates the search for optimal merging configurations using multi-objective black-box optimization algorithms, eliminating the reliance on human intuition and customized strategies.
(2) Enhanced Acquisition Strategy: To achieve high-quality model merging configurations with limited evaluation iterations, we introduced a weak-to-strong method to enhance acquisition function that allows previously evaluated superior configurations to guide new configurations’ searching. Besides, we used Fisher information to screen these configurations to increase the chance of discovering superior merging configurations.
(3) Additional Optimization Objective: In addition to task metrics, we have introduced a sparsity metric as an another optimization objective to enhance the model’s generalization performance across different tasks.
The remainder of this paper is organized as follows. Section 2 surveys the existing methods in model merging and multi-objective optimization. Section 3 introduces a novel multi-objective Bayesian optimization algorithm, MM-MO, tailored for model merging. In this section, we also present our improvements to the process of searching for new solutions and the design of new optimization objectives specifically for model merging scenarios. Section 4 presents the performance evaluation of the proposed method against related methods. Finally, we draw conclusions in Section 5.
2 Background
2.1 Model Merging
Generally speaking, there are four kinds of model merging (MM) methods.
One common approach is simply averaging the weights of different models, which has shown promising performance. Model soups [12] achieved notable performance improvements on computer vision tasks. Task arithmetic [13] merges LLMs by performing arithmetic operations on task vectors (also known as delta parameters), which represent the differences between the weights of fine-tuned LLMs and the original pre-trained model. Fisher-Weighted Averaging [14] is based on the Laplace approximation, where each model’s posterior is approximated as a Gaussian distribution, with its precision matrix corresponding to its Fisher information.
Spherical linear interpolation (SLERP) [15] can achieve smooth interpolation between two vectors and maintain directional consistency during model merging by calculating the angle between two vectors and determining the interpolation vector based on the interpolation factor and angle.
RegMean [16] merges different models by minimizing the prediction differences between the merged model and the source models.
To address the interference between parameters of different models caused by redundant parameter values and disagreements on the sign of parameters across models, TIES-Merging [17] introduces three novel steps when merging models: (1) resetting parameters with small delta parameters, (2) resolving sign conflicts, and (3) merging only the consistent parameters.
DARE [18] performs model merging by sparsifying delta parameters of source supervised fine-tuned (SFT) models with Drop and Re-Scale operations before actually merging them.
Recently, Daheim et al. [4] connected the inaccuracy of weighted-averaging to mismatches in the gradients and proposed a new uncertainty-based scheme to improve performance by reducing this mismatch.
Our approach builds upon the combination of TIES-Merging and DARE, leveraging their strengths to create a robust model merging strategy. By integrating the conflict resolution techniques of TIES-Merging with the parameter sparsification methods of DARE, we aim to automate the search for optimal model merging configurations, ensuring that the final merged model maintains high performance without the degradation typically associated with parameter conflicts.
2.2 Multi-objective Optimization
Without loss of generality, a multi-objective optimization problem can be stated as follows:
$&\min\textbf{ }\boldsymbol{f}(\boldsymbol{x})=(f_{1}(\boldsymbol{x}),f_{2}(\boldsymbol{x}),\ldots,f_{M}(\boldsymbol{x}))^{T}\\ &\textrm{ }s.t.\textrm{ }\boldsymbol{x}\in\mathcal{X}\end{aligned},$ (1)
where $\boldsymbol{x}$ $=$ $(x_{1},x_{2},\ldots,x_{d})$ is the decision vector, $\boldsymbol{f}(\cdot)$: $\mathcal{X}\rightarrow\mathcal{T}$ is $m$ objective functions, $\mathcal{X}$ and $\mathcal{T}$ denote the (nonempty) decision space and the objective space, respectively.
In order to compare the quality of solutions of an MOP, the concept of Pareto dominance is introduced:
Definition 2.1 (Pareto dominance [19])
Given two solutions $\boldsymbol{a}$, $\boldsymbol{b}\in$ $\mathcal{X}$, $\boldsymbol{a}$ is said to dominate $\boldsymbol{b}$ (denoted as $\boldsymbol{a}\prec\boldsymbol{b}$) if and only if $\forall i\in\{1,2,...,m\},f_{i}(\boldsymbol{a})\leq f_{i}(\boldsymbol{b})$ and $\exists j\in\{1,2,...,m\},f_{j}(\boldsymbol{a})<f_{j}(\boldsymbol{b})$. A solution $\boldsymbol{a}^{\ast}\in\mathcal{X}$ is Pareto optimal if no other solution $\boldsymbol{a}\in\mathcal{X}$ can dominate it. The solution set consisting of all the Pareto optimal solutions is called the Pareto set (PS): $PS$$=$$\{\boldsymbol{a}\in\mathcal{X}|\forall\boldsymbol{b}\in\mathcal{X},\boldsymbol{b}\not\prec\boldsymbol{a}\}$ and the corresponding objective vector set of the PS is the Pareto front (PF).
Multi-objective Optimization (MO) focuses on approximating the PS, targeting at a solution set with good convergence and diversity in the objective space. In this work, we utilize multi-objective optimization to address the complex challenge of model merging. By considering multiple objectives simultaneously, MO enables us to balance various performance metrics, such as different task-specific accuracies. This approach allows us to optimize the trade-offs between different performance criteria, leading to a more robust and effective merged model. Through this methodology, we aim to achieve a unified model that not only excels in specific tasks but also retains a broad and diverse set of capabilities from the individual models. This is particularly important in the context of LLMs, where different tasks may require different strengths and features from the base models. By leveraging MO, our method systematically explores the parameter space to identify optimal configurations, ensuring that the final model exhibits both high performance and comprehensive potential across various tasks.
3 Methodology
3.1 Overview
Figure 1: An illustration of automated model merging with multi-objective optimization (MM-MO).
In this section, we provide an overview of our proposed automated model merging approach, MM-MO. Our approach is designed to merge multiple pre-trained models into a single, unified model that retains the strengths of each individual model and surpasses the performance of any single model. Our automated approach aims at searching for the optimal merging configuration is discovered without relying on human intuition.
In Figure 1, we show the overall process of model merging based on multi-objective optimization, which is divided into three parts:
On the left is a schematic diagram of the automatic search process for model merging configurations. Our method starts with a set of source models, each obtained by performing SFT on different datasets using the same pre-trained backbone. These models have their own advantages for different task types. We merge these models to obtain a merged model, evaluate the performance of the merged model on different tasks as optimization objectives , and calculate the sparsity metric of the merged model as an additional optimization objective. For simplicity, we will denote $\boldsymbol{f}(\boldsymbol{x,Tj})$ later as $\boldsymbol{f_{j}}(\boldsymbol{x})$, $j\in\{1,...,M-1\}$. MM-MO generates new model merging configurations $\boldsymbol{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},...\}$ based on these objective values, allowing the creation of new merged models. This process is repeated until a preset evaluation iteration limit is reached, ultimately resulting in an optimal model merging configuration. In the middle is the complete process of using multi-objective Bayesian optimization to search for model merging configurations, detailed in Section 3.2. On the right is the step of generating new configurations. We enhanced the acquisition strategy of multi-objective Bayesian optimization by designing a weak-to-strong method to adjust candidate configurations, and we used Fisher information to select the next generation of model merging configurations, detailed in Section 3.3. We also present the pseudocode of the MM-MO framework in the appendix (Section A).
3.2 Model Merging via Multi-objective Optimization
In this subsection, we detail the multi-objective optimization process employed for model merging. Our method leverages parallel multi-objective bayesian optimization (qEHVI) [20]. This technique is an effective black-box optimization method designed to handle multiple objectives simultaneously, making it ideal for our model merging needs.
Our approach builds upon TIES-Merging with DARE [17, 21], utilizing qEHVI to search for the optimal model merging configuration. The integration of qEHVI allows for an automated and systematic exploration of the parameter space, identifying the best configurations that balance various performance metrics. Here are the detailed steps of our methodology.
1) Initialization: We start with a collection of pre-trained models, each fine-tuned on different tasks. These models and their configurations serve as the initial candidates for merging.
2) Objective Definition: We define multiple objectives that reflect the performance metrics we aim to optimize. Specifically, these objectives include task-specific accuracies, such as those from C-EVAL [22] and GSM8K [23], as well as a sparsity metric $\boldsymbol{g}(\boldsymbol{x})$ used in our experiments. These metrics are crucial as they ensure the merged model performs well across different types of tasks.
3) Surrogate Model Training: qEHVI leverages surrogate models to approximate the objective functions. We train these surrogate models on the initial set of model configurations and their corresponding performance metrics. This step involves creating a probabilistic model that predicts the performance of unseen configurations.
4) Acquisition Function Optimization: The key component of qEHVI is the Expected Hypervolume Improvement (EHVI) acquisition function. This function guides the search by quantifying the expected improvement in the objective space. We optimize this acquisition function to select the most promising model configurations to evaluate next. Specifically, the hypervolume (HV) is a measure of the space dominated by a set of solutions in the objective space, bounded from below by a reference point. The HV indicator is defined as follows:
$\textbf{HV}(S)=\Lambda(\{q\in\mathbb{R}^{d}|\exists p\in S:p\leq q\;\textbf{and}\;q\leq r\}),$ (2)
where $S$ signifies a solution set, $r$ is identified as a reference vector set, and $\Lambda(\cdot)$ denotes the Lebesgue measure. Hypervolume improvement (HVI) measures the additional hypervolume that a new point contributes beyond the existing Pareto frontier.
In black-box optimization, exact function values at unobserved points are unknown, making direct HVI computation infeasible. Instead, the EHVI acquisition function uses the surrogate model’s posterior distribution over the function values to compute the expected improvement. This is typically done using Monte Carlo (MC) integration to estimate EHVI [24], particularly when dealing with multiple candidates in parallel. The EHVI guides the search by identifying new points that are expected to provide the most significant improvement in the objective space. The more general parallel variant utilizing MC integration is given as follows:
$\begin{split}\alpha_{q\text{EHVI}}(\mathcal{X}_{\text{cand}}|\mathcal{P})&\approx\hat{\alpha}_{q\text{EHVI}}(\mathcal{X}_{\text{cand}}|\mathcal{P})\\ &=\frac{1}{N}\sum_{t=1}^{N}\text{HVI}(\tilde{\boldsymbol{f}}_{t}(\mathcal{X}_{\text{cand}})|\mathcal{P}),\end{split}$ (3)
where $\tilde{f}_{t}\sim p(\boldsymbol{f}|\mathcal{D})$ for $t=1,...,N$ and $\mathcal{X}_{\mathrm{cand}}=\{x_{i}\}_{i=1}^{q}$ [20]. After obtaining candidate configurations using the acquisition function, we adjusted the candidate configurations using the weak-to-strong method. We then selected the configurations based on their Fisher information to ultimately obtain a new generation of model merging configurations. For details, please refer to Section 3.3.
5) Parallel Evaluation: Unlike traditional methods, qEHVI supports parallel evaluations, which significantly speeds up the optimization process. In each iteration, we select a batch of model configurations and evaluate their performance on the defined objectives (e.g., C-EVAL and GSM8K accuracy).
6) Updating the Surrogate Model: After evaluating the selected configurations, we update the surrogate models with the new data. This iterative process continues, refining the surrogate models and improving the acquisition function’s accuracy in identifying high-potential configurations.
7) Convergence and Final Selection: The optimization process continues until convergence criteria are met, such as a maximum number of iterations or negligible improvement in the objective space. The final model merging configuration is selected based on the best trade-offs between the multiple performance metrics.
By using multi-objective Bayesian optimization method, our method systematically navigates the trade-offs between different task-specific accuracy, ensuring that the final merged model maintains high performance across various tasks. The iterative nature of this process allows for continuous improvement and fine-tuning, leading to a robust and high-performing unified model. This integration of qEHVI with TIES-Merging and DARE provides a powerful and automated solution for model merging, optimizing the balance between multiple performance criteria without relying on human intuition. This automated search for the optimal merging configuration ensures that the final model configuration surpasses the performance of any individual source model.
3.3 Enhanced Acquisition Strategy
To achieve high-quality model merging configurations within a limited number of evaluation iterations, we have further improved the acquisition function of multi-objective Bayesian optimization for the model merging scenario.
Specifically, we designed a weak-to-strong method, which adjusts the candidate configurations obtained using the acquisition function based on already evaluated superior configurations. In addition, we employ Fisher information to perform environment selection on the adjusted solutions, thereby further enhancing the chance of discovering superior model merging configurations.
The details of these two methods will be introduced in the following two sections.
3.3.1 Weak-to-strong Method
The Weak-to-Strong method has previously been utilized in aligning LLMs. EXPO [25] allows for the direct acquisition of a better-aligned model by extrapolating from the weights of an aligned model and its initial SFT checkpoint, eliminating the need for additional training. Inspired by this methodology, we developed a weak-to-strong method to improve the process of searching for new configurations in our model merging scenario.
The weak-to-strong method operates as follows:
1. Selection of Superior Configurations: From the already evaluated configurations, we select a small number (set to 5 in our experiments) of the top-performing configurations.
2. Differential Evolution: We apply Differential Evolution (DE) [26] to the parameters of these selected configurations. DE is a population-based optimization algorithm that creates new candidate solutions by combining the parameters of existing ones, ensuring diversity and exploration of the search space. Specifically, DE generates new candidate solutions using the following formula:
$\mathbf{v}_{i}^{(k+1)}=\mathbf{x}_{r_{1}}^{(k)}+F\cdot(\mathbf{x}_{r_{2}}^{(k)}-\mathbf{x}_{r_{3}}^{(k)})$ (4)
where $\mathbf{v}_{i}^{(k+1)}$ is the new solution vector for the $i$th individual at generation $k+1$. $\mathbf{x}_{r_{1}}^{(k)}$, $\mathbf{x}_{r_{2}}^{(k)}$, and $\mathbf{x}_{r_{3}}^{(k)}$ are randomly selected distinct individuals from the current population, and $r_{1}\neq r_{2}\neq r_{3}\neq i$. The differential mutation factor $F$ is a scaling factor typically between 0 and 1.
3. Stochastic Perturbation: To introduce an element of randomness and avoid local optima, we apply a stochastic perturbation with a probability of 50%. This perturbation replaces some of the parameters in the candidate configurations with those obtained from the DE process, encouraging exploration beyond the immediate neighborhood of the superior configurations.
This method helps in refining the acquisition function’s output by focusing the search around promising areas in the configuration space, thereby increasing the probability of finding high-quality model merging configurations.
3.3.2 Fisher Information
In the context of our model merging scenario, the purpose is to identify high-quality configurations within a limited number of evaluation iterations. To achieve this, we need to refine our surrogate model efficiently. One way to do this is by carefully selecting which configurations to evaluate based on their potential to improve the surrogate model’s accuracy. Fisher information [27] serves as a pivotal tool in this selection process. By prioritizing configurations with lower Fisher information for evaluation, we can enhance the precision of our surrogate model more efficiently.
Fisher information is a measure of the amount of information that an observable variable provides about an unknown parameter, thereby quantifying the expected information gain from data regarding parameter estimation. Mathematically, for a parameter $\theta$ and a probability density function $p(x;\theta)$ , the Fisher information $I(\theta)$ is expressed as:
$I(\theta)=-\mathbb{E}\left[\frac{\partial^{2}\log p(x;\theta)}{\partial\theta^{2}}\right]=\mathbb{E}\left[\left(\frac{\partial\log p(x;\theta)}{\partial\theta}\right)^{2}\right].$ (5)
In our methodology, the Fisher information for each candidate configuration is calculated using the current surrogate model, which is modeled as a Gaussian process [28], a commonly used surrogate model in Bayesian optimization. The detailed process is as follows:
1. Posterior Predictive Calculation: For each candidate configuration, the Gaussian process surrogate model is employed to compute the posterior predictive mean and variance.
2. Fisher Information Approximation: The variance of the posterior distribution is utilized to approximate the Fisher information for each configuration. Configurations with lower Fisher information indicate regions where the surrogate model’s predictions are less certain.
3. Environment Selection: Configurations with lower Fisher information are prioritized for evaluation. This ensures that configurations contributing more significantly to the model’s uncertainty are evaluated first, thereby enhancing the surrogate model’s overall accuracy.
By incorporating Fisher information into the environment selection process, we aim to improve the efficiency of the evaluation process, enhances the accuracy of the surrogate model more rapidly, and increases the possibility of discovering superior model merging configurations within the constraints of limited evaluation iterations.
3.4 Sparsity optimization objective
In addition to evaluating the performance of each merged model across various tasks, we introduce an additional optimization objective: a sparsity metric $\boldsymbol{g}(\boldsymbol{x})$ This metric is designed to quantify the sparsity of the delta parameters between the merged model and the base model. Specifically, $\boldsymbol{g}(\boldsymbol{x})$ is computed by taking the difference between the Merged Model (MM) and the Base Model to obtain the Delta Parameters, and then calculating the L1-Norm of all Delta Parameters. The sparsity metric $\boldsymbol{g}(\boldsymbol{x})$ can be expressed mathematically as:
$f_{M}(\boldsymbol{x})=\boldsymbol{g}(\boldsymbol{x})=\sum_{i=1}^{n}\|\Delta\theta_{i}\|_{1},$ (6)
where $\Delta\theta_{i}$ represents the delta parameters.
The purpose of incorporating this additional optimization objective is to prevent the final merged model from overfitting to specific evaluation tasks, thereby ensuring the merged model exhibits better overall performance and generalization capabilities.
The rationale behind designing this sparsity metric is supported by the findings of Yu et al. in their research on Dare [21], which demonstrated that for LLMs, removing up to 90% of the delta parameters does not significantly degrade performance. In some cases, the drop rate $p$ can even reach 99%, indicating that the supervised fine-tuning introduces substantial redundancy in the model parameters. This suggests that only minor modifications to the base model parameters are needed to unlock the model’s potential and achieve optimal results.
[…]
4 Experimental Setup
4.1 Source Models
To develop a model with strong comprehensive abilities across various disciplines, we applied our proposed model merging method to a set of source models, including Qwen1.5-7B-Chat [29], Liberated-Qwen1.5-7B [30], and firefly-qwen1.5-en-7B [31]. Besides these 7B models, we also experimented with models of different sizes and pre-trained backbones, including Qwen2-1.5B-Instruct [32], SauerkrautLM-1.5b [33], and Qwen2-1.5B-Ita [34]. Please see Table 5 in Section B.2 for their versions and correspondences with pre-trained backbones.
4.2 Datasets
For our study, we divided the usage of datasets into two distinct phases: configuration search and final model evaluation.
In the configuration search phase, we aimed to optimize the model parameters without overfitting to the test datasets. For this purpose, we used the C-EVAL [22] Validation Set, a Multi-Level Multi-Discipline Chinese Evaluation Suite, to ensure the model’s comprehensive abilities across different disciplines were appropriately calibrated. Additionally, to evaluate and optimize the mathematical reasoning abilities of our model, we sampled 1% of the GSM8K [23] Training Set. Since GSM8K does not have a predefined test set, the validation set is typically used for testing. To avoid data leakage, we selected a small, randomly chosen subset from the training set.
For the final evaluation of our models, we employed a range of datasets tailored to assess various capabilities, including C-EVAL [22], MMLU [35], GSM8K [23], HellaSwag [36], HumanEval [37], MBPP [38], WinoGrande [39] and BIG-Bench Hard (BBH) [40].
Among these, C-EVAL and MMLU are multidisciplinary examinations that evaluate the model’s comprehensive abilities; GSM8K assesses the model’s mathematical reasoning abilities; HellaSwag evaluates the model’s commonsense reasoning abilities; HumanEval and MBPP assess the model’s code generation abilities; WinoGrande evaluates the model’s language understanding abilities; and BBH assesses the model’s comprehensive reasoning abilities. Please see Section B.3 for more descriptions.
By utilizing these datasets, we ensured a comprehensive and balanced evaluation of models across a wide range of abilities and tasks.
4.3 Baselines
We have compared MM-MO with 5 state-of- the-art and classical merging methods, including Linear (Model Soups) [12], Task Arithmetic [13], TIES [17], DARE [21] and Model Breadcrumbs [41]. Since some model merging methods can be combined, there are a total of 7 baselines. Please see Section B.4 for more descriptions.
4.4 Evaluation Metrics
We calculate 5-shot accuracy for C-EVAL, 5-shot weighted accuracy for MMLU, 0-shot and 5-shot accuracy for GSM8K (0-shot for 7B model and 5-shot for smaller 1.5B model), 0-shot accuracy for HellaSwag, pass@1 for HumanEval and MBPP, 0-shot accuracy for WinoGrande, and 3-shot accuracy for BBH.
5 Experimental Results and Analysis
5.1 General Performance
Table 1: Performance Comparison of different merging methods and single models. (Model size: 7B params)
Merging Method Models Average Score C-EVAL GSM8K HellaSwag HumanEval MBPP MMLU WinoGrande
Single Model 1 Qwen1.5-7B-Chat 56.46 68.7 54.59 68.13 46.95 34.20 60.06 62.59
Single Model 2 Liberated-Qwen1.5-7B 57.29 69.7 53.30 71.02 48.78 38.80 58.84 60.62
Single Model 3 firefly-qwen1.5-en-7b 51.32 70.0 49.81 65.69 33.54 28.40 51.66 60.14
Linear (Model Soup) Single Model 1 + 2 + 3 58.88 71.1 54.89 72.72 50.61 39.80 60.80 62.27
Task Arithmetic Single Model 1 + 2 + 3 56.67 70.1 55.50 69.54 46.34 37.80 55.04 62.35
Dare + Task Arithmetic Single Model 1 + 2 + 3 58.38 69.8 55.27 69.97 51.22 39.40 60.23 62.75
Ties Single Model 1 + 2 + 3 53.32 65.7 53.15 69.58 34.76 29.00 58.05 62.98
DARE + Ties Single Model 1 + 2 + 3 57.78 69.5 55.72 69.23 49.39 37.80 60.06 62.75
Model Breadcrumbs Single Model 1 + 2 + 3 58.56 70.3 56.10 70.91 49.39 40.60 60.47 62.12
Model Breadcrumbs + Ties Single Model 1 + 2 + 3 58.41 70.2 55.72 70.59 50.00 39.80 60.35 62.19
MM-MO (Ours) Single Model 1 + 2 + 3 60.97 71.9 57.77 74.44 55.49 42.20 60.81 64.17
In the general experiment, we compared the performance of different model merging methods against individual models. The table presents the average scores and specific performances of each single model and their merged versions across various tasks.
Performance Analysis of Single Models. These three SFT models exhibited varying performances across the tasks. Specifically:
Single Model 1 performed best in the C-EVAL task (68.7) but had relatively lower scores in HumanEval (46.95) and MBPP (34.20).
Single Model 2 showed strong performance in HellaSwag (71.02) and MBPP (38.80) but underperformed in GSM8K (53.30) compared to Single Model 1.
Single Model 3 excelled in C-EVAL (70.0) but performed poorly in HumanEval (33.54) and MBPP (28.40).
These differences indicate that each single model has its own strengths and weaknesses across different tasks.
Comparison of Model Merging Strategies. Existing model merging methods surpassed individual models in certain tasks. For instance, the DARE + Ties method achieved higher scores in GSM8K (55.72) and HumanEval (49.39) than any single model. However, these methods do not guarantee significant improvements across all tasks. For example:
Linear merging performed well in C-EVAL (71.1) and HellaSwag (72.72), but its performance in WinoGrande (62.27) was lower than that of Single Model 1 and Single Model 3.
Task Arithmetic merging showed good results in GSM8K (55.50) but had an overall average score (56.67) lower than that of Single Model 2 (57.29).
DARE + Task Arithmetic improved scores in HumanEval (51.22) and MBPP (39.40) but did not surpass Single Models 1 and 2 in C-EVAL (69.8) and WinoGrande (62.75).
Ties method underperformed compared to individual models in some tasks, such as the average score (53.32) and HumanEval (34.76).
Breadcrumbs method generally performed well, with an average score of 58.56 and notable improvements in GSM8K (56.10) and MBPP (40.60), but did not perform as well as DARE + Task Arithmetic in HumanEval (49.39 vs. 51.22).
These results indicate that current methods do not ensure consistent improvements across all tasks. Existing model merging algorithms still have significant room for improvement in handling parameter conflicts and weight allocation during model merging. Therefore, we believe that further exploration of merging configurations can more effectively enhance the overall performance of merged models, achieving more comprehensive performance improvements.
Performance Analysis of Our method (MM-MO). MM-MO demonstrated the best overall performance, achieving the highest average score of 60.97 and it excelled in all evaluated tasks. This method’s superior performance across all tasks validates its effectiveness in optimizing model merging.
Furthermore, it is noteworthy that although we only used the accuracies of the C-EVAL and GSM8K datasets as optimization targets during the search configuration process, the final results show that the merged model obtained using MM-MO not only exhibits significant performance improvements on the C-EVAL and GSM8K test sets but also demonstrates noticeable enhancements in other types of tasks. This indicates that using MM-MO for model merging not only leverages the strengths of individual models in specific tasks but also enhances the overall potential of the model. It provides a reliable solution for integrating different single models into a high-performance and comprehensive model.
In summary, although the individual models show varying performance across different tasks, effective model merging strategies can substantially enhance overall performance. Our MM-MO method consistently achieved the best results in all evaluated tasks, validating the significant potential of merged models in tackling different tasks.
5.2 Comparison on BBH Dataset
Table 2: Average accuracy over 27 BBH tasks.
Merging Method Models BBH
Single Model 1 Qwen1.5-7B-Chat 38.02
Single Model 2 Liberated-Qwen1.5-7B 42.42
Single Model 3 firefly-qwen1.5-en-7b 28.41
MM-MO (Ours) Single Model 1 + 2 + 3 47.72
Figure 2: Performance of single and merged LLMs on different categories of the BBH dataset.
To validate our methods on difficult tasks, we utilize BBH, which includes a suite of 27 challenging BIG-Bench tasks that require multi-step reasoning. Table 2 shows the performance of single models and the MM-MO merged model on the BBH datasets. Our MM-MO merged model achieves an average BBH score of 47.72, significantly outperforming each of the single models. This result highlights the effectiveness of our model merging strategy in enhancing overall performance.
To further illustrate the advantages of our MM-MO method, we provide a detailed analysis of the performance across various sub-tasks within the BBH dataset. Figure 2 visualizes the scores of single models and the MM-MO merged model on different sub-tasks, As can be seen from the radar chart, MM-MO (red) covers the largest area and has better performance. Notably, in the Temporal Sequences task, MM-MO achieves a score of 47.60, outperforming the best individual model, which scored 40.80. In the Tracking Shuffled Objects (Three Objects) task, MM-MO reaches 54.40, significantly exceeding Single Model 2’s score of 42.00. Furthermore, in the Penguins in a Table task, MM-MO scores 62.33, surpassing Single Model 3’s 56.85. This demonstrates that MM-MO not only leverages the strengths of individual models but also explores their potential further.
5.3 Effectiveness of MM-MO with Different Parameter Scales and Source Models
Table 3: Performance comparison of different merging methods and single models. (Model size: 1.5B params)
Merging Method Models Average Score C-EVAL GSM8K HellaSwag HumanEval MBPP MMLU WinoGrande BBH
Single Model 4 Qwen2-1.5B-Instruct 47.05 68.2 62.09 40.77 42.68 30.00 54.58 42.70 35.37
Single Model 5 SauerkrautLM-1.5b 35.08 38.9 43.44 46.54 23.78 0.00 50.52 51.30 26.18
Single Model 6 Qwen2-1.5B-Ita 46.10 66.9 49.51 48.48 39.63 32.40 53.31 52.41 26.17
MM-MO (Ours) Single Model 4 + 5 + 6 51.04 69.3 63.15 50.65 43.90 35.60 54.84 54.54 36.33
To validate the versatility and generalization capabilities of our method, we applied MM-MO for model merging on source models with different parameter scales and different Pre-Trained Backbones, followed by extensive evaluations of the merged models. As shown in Table 3, the models obtained using MM-MO consistently outperformed any single model across all task types.
Furthermore, it is noteworthy that we selected three models that underwent SFT on different language datasets. Specifically, Qwen2-1.5B-Instruct was fine-tuned on Chinese and English datasets,
SauerkrautLM-1.5b on German, and Qwen2-1.5B-Ita on Italian. MM-MO demonstrated strong adaptability and generalization capabilities in merging models of different languages. This means that even models specializing in different languages can achieve significant performance improvements in multilingual tasks when merged using MM-MO. These results validate the broad applicability of the MM-MO method, showcasing its immense potential for merging models in multilingual, multi-task environments.
5.4 Ablation Study
Table 4: Ablation study results.
Merging Method Models Average Score C-EVAL C-EVAL (hard) GSM8K HumanEval
Single Model 1 Qwen1.5-7B-Chat 55.34 68.7 51.1 54.59 46.95
Single Model 2 Liberated-Qwen1.5-7B 55.47 69.7 50.1 53.30 48.78
Single Model 3 firefly-qwen1.5-en-7b 51.01 70.0 50.7 49.81 33.54
DARE + Ties Single Model 1 + 2 + 3 56.48 69.5 51.3 55.72 49.39
BO (C-EVAL) Single Model 1 + 2 + 3 58.62 71.4 51.5 56.71 54.88
BO (GSM8K) Single Model 1 + 2 + 3 58.01 70.3 51.9 57.39 52.44
MOBO Single Model 1 + 2 + 3 57.75 70.9 50.8 56.86 52.44
MOBO + Sparsity Single Model 1 + 2 + 3 58.56 71.5 52.6 57.09 53.05
MOBO + W2S Single Model 1 + 2 + 3 59.05 71.7 52.6 57.62 54.27
MOBO + Sparsity + W2S Single Model 1 + 2 + 3 58.83 72.0 52.1 57.54 53.66
MOBO + Sparsity + W2S + FI
MM-MO (Ours) Single Model 1 + 2 + 3 59.47 71.9 52.7 57.77 55.49
In this section, we conduct ablation studies on all modules of MM-MO to verify their respective effectiveness. We use the scores of the merged models obtained by the DARE-TIES without configuration search method on C-EVAL, GSM8K, and HumanEval as baselines.
1. BO (C-EVAL) & BO (GSM8K). First, we use a single-objective BO algorithm to search configurations for DARE-TIES model merging. As shown in Table 4 , BO (C-EVAL) achieves higher accuracy on C-EVAL, while BO (GSM8K) performs outstandingly on the GSM8K dataset. Additionally, HumanEval, which was not listed as an optimization target, also shows some improvement, indicating the importance and effectiveness of merging configuration searches for model merging.
2. MOBO. We employ multi-objective Bayesian optimization, considering both C-EVAL and GSM8K as targets. Compared to DARE + TIES, the model obtained using the MOBO method shows improvements in both C-EVAL and GSM8K. However, since no additional optimizations were made for the model merging scenario apart from adding more objectives, the overall improvement is limited.
3. MOBO + Sparsity. On the basis of MOBO, an additional optimization objective, the sparsity metric, is introduced. This metric helps prevent the final merged model from overfitting to the evaluated tasks, ensuring the merged model possesses stronger overall performance and better generalization. Experimental results also support this, showing that models obtained using this method perform better across all three types of tasks than those obtained using only MOBO.
4. MOBO + W2S. The acquisition strategy in MOBO is improved by introducing the weak-to-strong (W2S) method to adjust candidate configurations. This method focuses the search on promising regions of the configuration space, increasing the possibility of finding high-quality model merging configurations. This method also achieves better performance across all three types of tasks compared to using MOBO alone.
5. MOBO + Sparsity + W2S. On the basis of MOBO, additional sparsity metrics are introduced, and the weak-to-strong method is employed. The model merged using this method performs very well on C-EVAL, and also maintaining strong overall performance.
6. MOBO + Sparsity + W2S + Fisher. This is the MM-MO method proposed in this paper, which builds on MOBO by incorporating sparsity metrics, the weak-to-strong method, and Fisher information to select candidate solutions. These three contributions ensure that the final model merging configuration found has the best overall performance. As seen in Table 4, it achieved the highest average score of 59.47, with the highest accuracy on the C-EVAL (hard), GSM8K, and HumanEval.
This series of ablation studies demonstrates that the proposed improvements to the acquisition strategy (weak-to-strong method & Fisher information) and additional optimization objective (sparsity metric) all positively contribute to model merging configuration searches using MOBO. The weak-to-strong method focuses the search on promising areas within the configuration space, increasing the possibility of finding high-quality model merging configurations; incorporating Fisher information into the environment selection process can improve the accuracy of the surrogate model, increasing the chances of discovering excellent model merging configurations within a limited number of evaluations; the sparsity metric, as an additional optimization objective, helps prevent the final merged model from overfitting to the evaluated tasks, ensuring stronger overall performance and better generalization.
6 Limitation
Our experiments were limited by computational resources, restricting our method to merging models with up to 7 billion parameters. While results were promising within this range, the method’s effectiveness for larger models remains untested. Future research is needed to assess its scalability for more extensive models, which will be key to understanding the full potential of MM-MO.
Additionally, our method currently requires source models to be homologous (fine-tuned from the same pre-trained model) to ensure compatibility. This limits its use to scenarios with identical pre-trained models. Future work will explore extending our approach to merge heterogeneous models, expanding its applicability.
7 Conclusion
This paper proposes a novel approach (MM-MO) for LLM merging using multi-objective black-box optimization algorithms, addressing two major challenges: the reliance on human intuition for merging configurations and the difficulty of finding optimal configurations within a limited evaluation. Our contributions include automating the model merging process, enhancing model potential by multi-objective optimization. In addition, we improve the acquisition strategy of multi-objective optimization to deal with limited evaluations, and design an effective sparsity optimization objective to enhance the model’s generalization performance across different tasks. The experimental results demonstrate that our approach achieves superior performance compared to existing methods, offering a robust, high-performing solution for integrating diverse models.
Appendix A Pseudocode of MM-MO
Input :  number of iterations $K$, batch size $B$, number of initial configurations $N$
Output : final configuration $\boldsymbol{X}_{Final}$
1 Initialize $N$ configurations $\boldsymbol{X}_{0}$ randomly
2 Evaluate $\boldsymbol{Y}_{0}$ using objectives $\boldsymbol{f_{1}}(\boldsymbol{X}_{0}),\boldsymbol{f_{2}}(\boldsymbol{X}_{0}),...,\boldsymbol{g}(\boldsymbol{X}_{0})$
3 for $k=0$ to $K-1$ do
4       Train surrogate model $GP_{i}^{k}$ based on $\{\boldsymbol{X}_{k},\boldsymbol{Y}_{k}\}$ for each objectives.
5       ${\boldsymbol{X}}_{k}^{B}\leftarrow$ Generate new configurations based on $GP^{k}$ and $\{\boldsymbol{X}_{k},\boldsymbol{Y}_{k}\}$ (Algorithm 2)
6       Update $\boldsymbol{X}_{k+1}\leftarrow\boldsymbol{X}_{k}\cup{\boldsymbol{X}_{k}^{B}}$
7       Evaluate $\boldsymbol{Y}_{k+1}$ using $\boldsymbol{f_{1}}(\boldsymbol{X}_{k}^{B}),\boldsymbol{f_{2}}(\boldsymbol{X}_{k}^{B}),...,\boldsymbol{g}(\boldsymbol{X}_{k}^{B})$
8 end for
$\boldsymbol{X}_{Final}\leftarrow$ Final selection from $\boldsymbol{X}_{K}$
Algorithm 1 The framework of MM-MO
Input : evaluated configurations $\{\boldsymbol{X}_{E},\boldsymbol{Y}_{E}\}$, acquisition function $\boldsymbol{qEHVI}$, surrogate model $GP$
Output : next-generation configuration $\boldsymbol{X}_{New}$
1 $\boldsymbol{X}_{C}\leftarrow$ Generate candidate configurations based on $\boldsymbol{qEHVI}$ and $\{\boldsymbol{X}_{E},\boldsymbol{Y}_{E}\}$
2 $\boldsymbol{X}_{H}\leftarrow$ Select high-quality configurations based on $\{\boldsymbol{X}_{E},\boldsymbol{Y}_{E}\}$
3 $\boldsymbol{X}_{A}\leftarrow$ Adjust configurations by weak-to-strong method based on $\boldsymbol{X}_{C}$ and $\boldsymbol{X}_{H}$
$\boldsymbol{X}_{New}\leftarrow$ Select the next-generation configuration by Fisher information based on $\boldsymbol{X}_{A}$
Algorithm 2 Generate new configuration
In this section, we present the pseudocode for the MM-MO framework (Algorithm 1) and for generating new configurations (Algorithm 2). The basic steps of the overall MM-MO process are as follows:
1. Initialization Phase.
The process begins by randomly generating an initial set of $N$ configurations, denoted as $\boldsymbol{X}_{0}$, where $N$ defaults to 10.
These configurations are evaluated using the evaluation function $\boldsymbol{f}(\boldsymbol{x})$ and the sparsity metric calculating function $\boldsymbol{g}(\boldsymbol{x})$. The results are combined into an initial set of objective values, $\boldsymbol{Y}_{0}$, which includes both evaluation results and sparsity metrics.
2. Iterative Optimization Phase:
The algorithm iterates $K$ times, where $K$ defaults to 4. In each iteration, the current set of configurations $\boldsymbol{X}_{k}$ and their associated objective values $\boldsymbol{Y}_{k}$ are used to train surrogate models (Gaussian Processes, GP) for each objective, which is a commonly used surrogate model in Bayesian optimization.
Based on these surrogate models and the current evaluated configurations, a batch of new candidate configurations ${\boldsymbol{X}}_{k}^{B}$ is generated. We set the batch size to 5 in our experiment.
The newly generated configurations are added to the existing set, and their objective and sparsity metrics are evaluated. This updates both the configuration set and their objective values for the next iteration.
3. Final Selection Phase:
After completing all iterations, the configuration with the highest objective values is selected as the final configuration $\boldsymbol{X}_{Final}$, which serves as the output of the algorithm.
During the generation of new configurations, the algorithm first generates candidate configurations based on the acquisition function (e.g., qEHVI [20]) and the previously evaluated configurations. It then selects high-quality configurations based on objective values, applies a weak-to-strong adjustment method, and finally selects half of the configurations according to the Fisher information for the next generation. This approach can increase the chances of discovering excellent model merging configurations within a limited number of evaluations. For details, please refer to Section 3.3.
Appendix B Detailed Experimental Settings
B.1 Implementation Details
The maximal number of generated token is set to 1024 on all datasets. Boolean parameter do_sample is set to False for greedy decoding (always choosing the most probable next token). Experiments are conducted on one NVIDIA A100 GPU.
B.2 Details of SFT and Pre-Trained Backbones of LMs
Table 5 shows the versions and correspondences with pre-trained backbones of SFT LMs.
Table 5: Versions and correspondences with pre-trained backbones of SFT LMs.
Parameter Scale SFT LMs URL Pre-Trained Backbones
7B Qwen1.5-7B-Chat [29] https://huggingface.co/Qwen/Qwen1.5-7B-Chat Qwen1.5-7B [29]
Liberated-Qwen1.5-7B [30] https://huggingface.co/abacusai/Liberated-Qwen1.5-7B
firefly-qwen1.5-en-7B [31] https://huggingface.co/YeungNLP/firefly-qwen1.5-en-7b
1.5B Qwen2-1.5B-Instruct [32] https://huggingface.co/Qwen/Qwen2-1.5B-Instruct Qwen2-1.5B [32]
SauerkrautLM-1.5b [33] https://huggingface.co/VAGOsolutions/SauerkrautLM-1.5b
Qwen2-1.5B-Ita [34] https://huggingface.co/DeepMount00/Qwen2-1.5B-Ita
B.3 Descriptions of Datasets
In this section, we describe the datasets used in our study, each serving a specific purpose in evaluating various model abilities.
  • C-EVAL [22] and MMLU [35]. C-EVAL and MMLU are multidisciplinary examinations assessing a model’s general knowledge across various subjects such as humanities, sciences, and social studies.
  • GSM8K [23]. The GSM8K dataset focuses on evaluating mathematical reasoning abilities through a collection of grade school math problems, testing the model’s problem-solving skills in mathematics.
  • HellaSwag [36]. HellaSwag assesses commonsense reasoning by presenting multiple-choice questions that require the model to complete scenarios based on everyday common sense.
  • HumanEval [37] and MBPP [38]. HumanEval and MBPP are used to assess code generation abilities. HumanEval contains code snippets and problem descriptions, while MBPP includes programming problems to test code generation accuracy.
  • WinoGrande [39]. WinoGrande evaluates language understanding by presenting sentences with ambiguous pronouns, requiring the model to identify the correct antecedents.
  • BIG-Bench Hard (BBH) [40]. BBH assesses comprehensive reasoning abilities with challenging questions that test advanced reasoning and problem-solving skills.
B.4 Descriptions of Model Merging Methods
We experiment with five model merging methods:
  • Linear (Model Soups) [12]. Model Soups improves model accuracy by averaging the weights of multiple models fine-tuned with different hyperparameters. This method enhances robustness without additional inference or memory costs, making it effective for large pre-trained models in tasks like image classification and natural language processing.
  • Task Arithmetic [13]. Task Arithmetic uses task vectors, which represent directions in the weight space of a pre-trained model, to modify and combine model behaviors. By performing arithmetic operations on these vectors, the model’s performance can be steered across multiple tasks efficiently and effectively, without requiring additional training data for each task.
  • TIES [17]. TIES (TRIM, ELECT SIGN & MERGE) addresses parameter interference in model merging by resetting minimally changed parameters, resolving sign conflicts, and merging only aligned parameters. This method outperforms existing techniques in various settings, maintaining valuable information and ensuring better performance in multi-task models.
  • DARE [21]. DARE (Drop And Rescale) sparsifies delta parameters by setting most to zero and rescaling the rest. This approach mitigates parameter interference when merging multiple fine-tuned models into one, enhancing capabilities without retraining. DARE proves particularly effective for large-scale language models, often surpassing the performance of individual source models.
  • Model Breadcrumbs [41]. Model Breadcrumbs enhance task performance by defining a sparse set of weights that create a trajectory within the weight space of a pre-trained model. This method improves efficiency and performance across multiple tasks without needing hyperparameter tuning.
Appendix C Additional Experiments
C.1 Comparison on C-EVAL Dataset
Figure 3: Performance of Single Model and Merged Model (MM-MO) on C-EVAL.
Figure 3 visualizes the accuracy of single models and the MM-MO merged model in answering four different categories of C-EVAL questions, the average C-EVAL accuracy, and the accuracy on the hard C-EVAL questions. As shown in the figure, MM-MO (red) covers the largest area and demonstrates superior performance. Additionally, in all four categories of C-EVAL tasks—STEM, Social Sciences, Humanities, and Others—the MM-MO merged model achieves higher accuracy than any single best-performing model, which intuitively demonstrates that using MM-MO for model merging can combine the strengths of various models and unlock their potential.
C.2 Case Study: Mathematical Reasoning
Table 6: Case Study on GSM8K.
Problem
Instruction Mark took a test yesterday that consisted of 75 questions. He completed the test at a rate of 5 questions per hour. Today, he took another test of 100 questions at the same rate. If Mark had 8 hours to complete the first test and 6 hours to complete the second one, how many questions did he leave incomplete?
Ground Truth Answer 105
Note Questions Mark completed in yesterday’s test: 5*8=40. Questions Mark completed in today’s test: 5*6=30. Questions Mark left incomplete in yesterday’s test: 75-40=35 Questions Mark left incomplete in today’s test: 100-30=70 Missed questions in total 35+70=105.
Answers
MM-MO (Ours) To find out how many questions Mark left incomplete, we need to calculate the number of questions he completed in each test and then subtract that from the total number of questions in each test. For the first test: - Mark completed 5 questions per hour for 8 hours, so he completed 5 * 8 = 40 questions. - The total number of questions in the first test is 75, so he left 75 - 40 = 35 questions incomplete. For the second test: - Mark completed 5 questions per hour for 6 hours, so he completed 5 * 6 = 30 questions. - The total number of questions in the second test is 100, so he left 100 - 30 = 70 questions incomplete. Now, to find the total number of questions left incomplete, we add the number of questions left incomplete in each test: 35 (first test) + 70 (second test) = 105 questions The answer is: 105 questions.
TIES To find out how many questions Mark left incomplete, we need to calculate the total number of questions he completed in each test and then subtract that from the total number of questions in each test. For the first test: - Number of questions completed: 75 questions / 5 questions per hour = 15 hours - Number of questions left incomplete: 8 hours - 15 hours = -7 hours (This is not possible, so there are no questions left incomplete for the first test) For the second test: - Number of questions completed: 100 questions / 5 questions per hour = 20 hours - Number of questions left incomplete: 6 hours - 20 hours = -14 hours (This is not possible, so there are no questions left incomplete for the second test)
DARE-TIES You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day! You too! Thank you! You’re welcome! Have a great day!
Table 6 provides a comparative analysis of responses to a mathematical reasoning problem by three models. By employing MM-MO to search for the optimal configuration for model merging, we enhanced our model’s performance in mathematical reasoning capabilities.
Specifically, in this problem, only our merged model gives the correct answer. Our model (merged by MM-MO method) understood the question well and followed the step by step instructions to give the final correct answer. However, the model merged by TIES method, while also following the step by step instructions, does not properly understand the problem and give a mathematical reasoning process that meets our requirements. Finally, the model merged by DARE-TIES method fails entirely, providing irrelevant and repetitive responses, despite DARE-TIES utilizing a dropout-like method to avoid parameter conflicts during model merging, which generally results in better performance compared to TIES, it can suffer from significant performance drops on specific problems if the model merging configuration is not correctly set. This issue is evident in the provided case study, where DARE-TIES fails to deliver a coherent response. This highlights the importance of correctly configuring model merges to retain crucial parameters, ensuring consistent performance across diverse problem sets. The MM-MO model’s success demonstrates the effectiveness of our model merging strategy in overcoming these challenges and providing accurate solutions.
C.3 Parameter Sensitivity Experiment
Table 7: Performance of MM-MO with different number of function evaluations.
Merging Method Models Average Score C-EVAL C-EVAL (hard) GSM8K HumanEval
Single Model 1 Qwen1.5-7B-Chat 55.34 68.7 51.1 54.59 46.95
Single Model 2 Liberated-Qwen1.5-7B 55.47 69.7 50.1 53.30 48.78
Single Model 3 firefly-qwen1.5-en-7b 51.01 70.0 50.7 49.81 33.54
MM-MO / FE = 20 Single Model 1 + 2 + 3 54.62 70.7 51.5 53.60 42.68
MM-MO / FE = 40 Single Model 1 + 2 + 3 59.27 72.0 52.8 57.62 54.66
MM-MO / FE = 30 (Default) Single Model 1 + 2 + 3 59.47 71.9 52.7 57.77 55.49
In this section, we compare the performance of MM-MO with different number of function evaluations to validate the sensitivity of our proposed method to this parameter. Table 7 illustrates the impact of varying number of function evaluations on the performance of MM-MO.
When FE=20, the performance of the merged model across various tasks is consistently inferior to that of FE=30 and FE=40. This suggests that a function evaluation count of 20 is insufficient to discover an optimal model merging configuration. Comparing the performances of the merged model between FE=30 and FE=40, we observe that while the accuracy on the C-EVAL under FE=30 is lower than under FE=40, its performance on GSM8K and HumanEval slightly surpasses that of the latter. This indicates that a function evaluation count of 30 is sufficient for MM-MO to search for a satisfactory model merging configuration. Therefore, considering the trade-off between computational cost and performance, we set the default number of function evaluations in this study to 30.
C.4 Time Overhead of MM-MO
Table 8: Time overhead of MM-MO for different model size.
Merging Method Model Size Time Overhead (min)
MM-MO 1.5B 259.21
MM-MO 7B 526.52
Table 8 lists the time overhead of MM-MO for different model size, 1.5B and 7B. Compared to fine-tuning on large datasets to improve the overall performance of the model, using MM-MO for model merging with existing models does not require additional training data. It only needs to be evaluated on the existing language model evaluation datasets to quickly obtain a merged model with great overall performance in a relatively short period of time.
C.5 The Final Configuration Searched using MM-MO
Figure 4: Final Configuration Searched using MM-MO. (1.5B & 7B)
Figure 4 illustrates the final parameter configuration of model merging (1.5B & 7B) searched using MM-MO. Density is the fraction of weights in differences from the base model to retain and weight represents the relative weighting of a given model. From the figure we can know the contribution of different single models to the final model merging.
Appendix D Licenses
The licenses for the codes we used in this work are shown in Table 9.
Table 9: List of licenses for the codes suite we used in this work.
Resources Type Link License
MergeKit [42] Code https://github.com/arcee-ai/mergekit LGPL-3.0 license
OpenCompass [43] Code https://github.com/open-compass/opencompass Apache-2.0 license